OpenStudy (anonymous):

Simplify (8z - 10) ÷ (-2) + 5(z - 1). z z - 1 z - 2 I got b

OpenStudy (anonymous):

i got z-1.5

jigglypuff314 (jigglypuff314):

÷ (-2) is like * (-1/2) so (8z - 10) * (-1/2) + 5(z - 1) then use the order of operations PEMDAS (-1/2)(8z) + (-1/2)(-10) + 5z + (5)(-1) then simplify -4z + 5 + 5z - 5 then add/subtract like terms

OpenStudy (anonymous):

so am i correct?

jigglypuff314 (jigglypuff314):

not quite :) try adding like terms for -4z + 5 + 5z - 5

OpenStudy (anonymous):

yeh so x-1

OpenStudy (anonymous):

z*

jigglypuff314 (jigglypuff314):

-4z + 5z + 5 - 5 z ^0 so it would just be z

OpenStudy (anonymous):

isnt -4 +5=1

jigglypuff314 (jigglypuff314):

-4z + 5z = z(-4 + 5) = (1)z = z 5 - 5 = 0